# Environment Maps: Structured Environmental Representations for Long-Horizon Agents

### 26 Mar 2026 v2

@@ 1 Introduction
To bridge this gap, we introduce extitEnvironment Maps, a persistent, agent-agnostic representation that consolidates heterogeneous experience (e.g., screen recordings, execution traces, structured workflows, documentation) into a queryable graph by segmenting workflows, clustering page contexts, and parameterizing actions e.g., “Click John Smith” $\rightarrow$ “Click {customer_name}”). Environment maps encode abstract contexts, parameterized actions, observed workflows, and tacit domain knowledge, separating durable environment knowledge from UI details and task-specific policies to support reuse across sessions and UI versions. In our experiments, agents equipped with environment maps achieve 28.2% success on WebArena, nearly doubling the 14.2% baseline rate of a extttclaude-sonnet-4-5 agent without map access, and outperform agents with access to the raw trajectories used to generate the maps (23.3%), demonstrating that representation plays a significant role beyond raw data itself. Our contributions are as follows:

### 27 Mar 2026 v3

@@ 1 Introduction
To bridge this gap, we introduce Environment Maps, a persistent, agent-agnostic representation that consolidates heterogeneous experience (e.g., screen recordings, execution traces, structured workflows, documentation) into a queryable graph by segmenting workflows, clustering page contexts, and parameterizing actions e.g., “Click John Smith” $\rightarrow$ “Click {customer_name}”). Environment maps encode abstract contexts, parameterized actions, observed workflows, and tacit domain knowledge, separating durable environment knowledge from UI details and task-specific policies to support reuse across sessions and UI versions. In our experiments, agents equipped with environment maps achieve 28.2% success on WebArena, nearly doubling the 14.2% baseline rate of a claude-sonnet-4-5 agent without map access, and outperform agents with access to the raw trajectories used to generate the maps (23.3%), demonstrating that representation plays a significant role beyond raw data itself. Our contributions are as follows: